# On the complexity of matrix Putinar’s Positivstellensatz

### v1

@@ 5. Degree bounds for matrix Positivstellensatze
5.1. Proof of Theorem 1.1
For the polynomial matrix $G(x)$, let $d_{1},\dots,d_{\theta(m)}$ be polynomials as in Theorem 4.2. Note that we assume that
@@
(5.2) $\mathcal{G}(x)=-\min\left(\frac{d_{1}(x)}{\left\|d_{1}(x)\right\|_{B}},\ldots,\frac{d_{\theta(m)}(x)}{\left\|d_{\theta(m)}(x)\right\|_{B}},0\right).$
If $\mathcal{G}(x)=0$, we have that $x\in K$, i.e., $d(x,K)=0$. By Corollary 2.2, there exist positive constants $\eta$, $\kappa$ such that the following Lojasiewicz inequality holds:
@@
In the following, we prove Theorem 1.1.
@@
Let $K$ be as in (5.1) with $m\geq 2$. Suppose that $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix with $\deg(F)=d$, and $(1-\|x\|_{2}^{2})\cdot I_{\ell}\in\mbox{QM}[G]^{\ell}$. If $F(x)\succeq F_{\min}\cdot I_{m}\succ 0$ on $K$, then $F\in\mbox{QM}[G]_{k}^{\ell}$ for
@@
In the above, $\kappa,\eta$ are respectively Lojasiewicz constant and exponent as in (5.3) and $C>0$ is a constant independent of $F$, $G$.
Proof.
@@
$K=\{d_{1}(x)\geq 0,\dots,d_{\theta(m)}(x)\geq 0\}.$
It follows from Theorem 3.3 that for
$k\geq C\cdot{8}^{7\eta}\cdot 3^{6(m-1)}\theta(m)^{3}n^{2}d_{G}^{6}\kappa^{7}d^{14\eta}\left(\frac{\|F\|_{B}}{F_{\min}}\right)^{7\eta+3},$
there exist matrices $P_{\alpha}^{(k)}\succ 0$, SOS polynomial matrices $S_{i}\in\Sigma[x]_{k}^{\ell}$ ($i=0,\dots,\theta(m)+1$) with $\deg(S_{0})\leq k$, $\deg(d_{i}S_{i})\leq k~{}(i=1,\dots,\theta(m))$ such that
$F=S_{0}+d_{1}S_{1}+\cdots d_{\theta(m)}S_{\theta(m)}+\sum_{\alpha\in\mathbb{N}_{k}^{n}}P_{\alpha}^{(k)}B_{\alpha}^{k}(x).$
Since $(1-\|x\|_{2}^{2})\cdot I_{\ell}\in\mbox{QM}[G]^{\ell}$, we know that $1-\|x\|_{2}^{2}\in\mbox{QM}[G]$. Note that $d_{1},d_{2},\dots,d_{\theta(m)}\in QM[G]$ and $1+x_{1},\cdots,1+x_{n},\sqrt{n}-x_{1}-\cdots-x_{n}\in\mbox{QM}[1-\|x\|_{2}^{2}]$. The proof is completed. ∎∎
@@
Theorem 5.1 implies that for $k$ greater than the quantity as in (5.4), we have $F\in\mbox{QM}[G]_{k}^{\ell}$. However, the bound (3.6) provided in Theorem 3.3 is typically smaller than that as in (5.4), and the proof of Theorem 5.1 builds up on Theorem 3.3.
5.2. Proof of Theorem 1.2
For a polynomial matrix $F=\sum\limits_{\alpha\in\mathbb{N}^{n}_{d}}F_{\alpha}x^{\alpha}$ with $\deg(F)=d$, $F^{hom}$ denotes its highest-degree homogeneous terms, i.e.,
$F^{hom}(x)=\sum\limits_{\alpha\in\mathbb{N}^{n},|\alpha|=d}F_{\alpha}x^{\alpha}.$
The homogenization of $F$ is denoted by $\widetilde{F}$, i.e.,
$\widetilde{F}(\tilde{x})=\sum_{\alpha\in\mathbb{N}^{n}_{d}}F_{\alpha}x^{\alpha}x_{0}^{d-|\alpha|}$
for $\tilde{x}:=(x_{0},x)$. Let
@@
Proof.
Suppose $\tilde{x}=(x_{0},x)\in\widetilde{K}$. If $x_{0}=0$, we have that
$\widetilde{F}(\tilde{x})=F^{hom}(x)\succ 0.$
If $x_{0}\neq 0$, then
@@
which implies that $x/x_{0}\in K$. Since $F^{hom}\succ 0$ on $\mathbb{R}^{n}$, the degree of $F$ must be even and we know that
@@
Since $\widetilde{K}$ is compact, the optimal value of (5.6) is positive, i.e., $\widetilde{F}_{\min}>0$. ∎∎
For the polynomial matrix $(x_{0})^{d_{0}}\widetilde{G}(\tilde{x})$, let $d_{1}(\tilde{x}),\dots,d_{\theta(m)}(\tilde{x})$ be the scalar polynomials as in Theorem 4.2. Denote
(5.7) $\mathcal{G}(\tilde{x})=-\min\left(\frac{d_{1}(\tilde{x})}{\left\|d_{1}(\tilde{x})\right\|_{B}},\ldots,\frac{d_{\theta(m)}(\tilde{x})}{\left\|d_{\theta(m)}(\tilde{x})\right\|_{B}},\frac{1-\|\tilde{x}\|_{2}^{2}}{\|1-\|\tilde{x}\|_{2}^{2}\|_{B}},\frac{\|\tilde{x}\|_{2}^{2}-1}{\|1-\|\tilde{x}\|_{2}^{2}\|_{B}},0\right).$
@@
In the following, we prove Theorem 1.2. It provides a polynomial bound on the complexity of matrix Putinar–Vasilescu’s Positivstellensatz.
@@
In the above, $\kappa$, $\eta$ are respectively Lojasiewicz constant and exponent as in (5.8), $C>0$ is a constant independent of $F$, $G$ and $\widetilde{F}_{\min}$ is the optimal value of (5.6).
Proof.
By Lemma 5.2, we have that $\widetilde{F}\succeq\widetilde{F}_{\min}\cdot I_{\ell}\succ 0$ on $\widetilde{K}$. Note that $\mbox{QM}[\|\tilde{x}\|^{2}-1]^{\ell}$ is archimedean and $K$ can be equivalently described as
$d_{1}(\tilde{x})\geq 0,\dots,d_{\theta(m)}(\tilde{x})\geq 0,1-\|\tilde{x}\|_{2}^{2}\geq 0,\|\tilde{x}\|_{2}^{2}-1\geq 0,$
@@
Suppose that $S=V^{T}V$ for some polynomial matrix $V$. Then, we know that there exist polynomial matrices $V^{(1)},V^{(2)}$ with $2\deg(V^{(1)})\leq\deg(S)$, $2\deg(V^{(2)})+2\leq\deg(S)$ such that
@@
Then, it holds that
$\begin{array}[]{ll}S\left(\frac{(1,x)}{\sqrt{1+\|x\|_{2}^{2}}}\right)=&\frac{(V^{(1)})^{T}V^{(1)}+(1+\|x\|_{2}^{2})(V^{(2)})^{T}V^{(2)}}{\sqrt{1+\|x\|_{2}^{2}}^{\deg(S)}}\\ &+\frac{(V^{(2)})^{T}V^{(1)}+(V^{(1)})^{T}V^{(2)}}{\sqrt{1+\|x\|_{2}^{2}}^{\deg(S)-1}}.\end{array}$
Similarly, there exist polynomial matrices $P_{i}^{(1)},P_{i}^{(2)}$ with $\deg(P_{i}^{(1)})\leq\deg(P_{i})$, $\deg(P_{i}^{(2)})+1\leq\deg(P_{i})$ such that
@@
$\begin{array}[]{l}\sum\limits_{i=1}^{t}P_{i}\left(\frac{(1,x)}{\sqrt{1+\|x\|_{2}^{2}}}\right)^{T}\frac{G(x)}{\sqrt{1+\|x\|_{2}^{2}}^{d_{0}+d_{G}}}P_{i}\left(\frac{(1,x)}{\sqrt{1+\|x\|_{2}^{2}}}\right)\\ \quad\quad\quad\quad\quad\quad=\sum_{i=1}^{t}\frac{(P_{i}^{(1)})^{T}G(x)P_{i}^{(1)}+(1+\|x\|_{2}^{2})(P_{i}^{(2)})^{T}G(x)P_{i}^{(2)}}{\sqrt{1+\|x\|_{2}^{2}}^{d_{0}+d_{G}+2\deg(P_{i})}}\\ \quad\quad\quad\quad\quad\quad\,+\frac{(P_{i}^{(2)})^{T}G(x)P_{i}^{(1)}+(P_{i}^{(1)})^{T}G(x)P_{i}^{(2)}}{\sqrt{1+\|x\|_{2}^{2}}^{d_{0}+d_{G}+2\deg(P_{i})-1}}.\\ \end{array}$
The identity (5.10) implies that for $k$ satisfying (5.9), we have
@@
for some $Q\in\mbox{QM}[G]^{\ell}$ with $\deg(Q)\leq 2k+d$ and symmetric polynomial matrix $H$. Since $\sqrt{1+\|x\|_{2}^{2}}$ is not a rational function, we have that $(1+\|x\|_{2}^{2})^{k}F=Q$ and thus $(1+\|x\|_{2}^{2})^{k}F\in\mbox{QM}[G]^{\ell}_{2k+d}$. ∎∎
Corollary 5.4.
Let $K$ be as in (5.1) with $m\geq 2$. Suppose that $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix with $\deg(F)=d$. If $F\succeq 0$ on $K$, then for $\varepsilon>0$, we have
if $k\geq C\cdot\varepsilon^{-7\eta-3}.$ Here, $\eta$ is the Lojasiewicz exponent as in (5.8) and $C>0$ is a constant depending on $F$, $G$.
Proof.
Note that $F\succeq 0$ on $K$ and
$d+2\geq\deg(F+\varepsilon\cdot(1+\|x\|_{2}^{2})^{\lceil\frac{d+1}{2}\rceil})=2\lceil\frac{d+1}{2}\rceil>d.$
Hence, we know that for $\tilde{x}\in\widetilde{K}$, the homogenization of $F+\varepsilon\cdot(1+\|x\|_{2}^{2})^{\lceil\frac{d+1}{2}\rceil}I_{\ell}$ satisfies
$\begin{array}[]{ll}x_{0}^{2\lceil\frac{d+1}{2}\rceil-d}\widetilde{F}+\varepsilon\cdot(x_{0}^{2}+\|x\|_{2}^{2})^{\lceil\frac{d+1}{2}\rceil}I_{\ell}&\succeq\varepsilon\cdot(x_{0}^{2}+\|x\|_{2}^{2})^{\lceil\frac{d+1}{2}\rceil}I_{\ell}\\ &\succeq\varepsilon\cdot I_{\ell}.\\ \end{array}$
The conclusion follows from Theorem 5.3. ∎∎
5.3. Convergence rate of matrix SOS relaxations

### v2

@@ 5 Degree bounds for matrix Positivstellensätze
5.1 Proof of Theorem 1.1
For the polynomial matrix $G(x)$, let $d_{1},\dots,d_{\theta(m)}$ be the polynomials as in Theorem 4.1. Note that we assume
@@
(5.2) $\mathcal{G}(x)=-\min\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{d_{1}(x)}{\left\|d_{1}\right\|_{B}}},\ldots,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{d_{\theta(m)}(x)}{\left\|d_{\theta(m)}\right\|_{B}}},0\right).$
If $\mathcal{G}(x)=0$, we have that $x\in K$, i.e., $d(x,K)=0$. By Corollary 2.2, there exist positive constants $\eta$, $\kappa$ such that the following Łojasiewicz inequality holds:
@@
In the following, we give the proof of Theorem 1.1.
@@
Let $K$ be as in (5.1) with $m\geq 2$. Suppose that $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix with $\deg(F)=d$, and $(1-\|x\|_{2}^{2})\cdot I_{\ell}\in\mbox{QM}[G]^{\ell}$. If $F(x)\succeq F_{\min}\cdot I_{\ell}\succ 0$ on $K$, then $F\in\mbox{QM}[G]_{k}^{\ell}$ for
@@
In the above, $\kappa,\eta$ are respectively the Łojasiewicz constant and exponent in (5.3) and $C>0$ is a constant independent of $F$, $G$. Furthermore, the Łojasiewicz exponent can be estimated as
$\eta\leq(3^{m-1}d_{G}+1)(3^{m}d_{G})^{n+\theta(m)-2}.$
Proof 5.2.
@@
$K=\{x\in\mathbb{R}^{n}\mid d_{1}(x)\geq 0,\dots,d_{\theta(m)}(x)\geq 0\}.$
It follows from Theorem 3.5 that $F\in QM[D]_{k}^{\ell}$, where $D$ is the diagonal matrix with diagonal entries $d_{1},\ldots,d_{\theta(m)}$, $1-\|x\|_{2}^{2}$, for
$k\geq C\cdot{8}^{7\eta}\cdot 3^{6(m-1)}\theta(m)^{3}n^{2}d_{G}^{6}\kappa^{7}d^{14\eta}\left(\frac{\|F\|_{B}}{F_{\min}}\right)^{7\eta+3}.$
Equivalently, there exist SOS polynomial matrices $S_{i}\in\Sigma[x]^{\ell}$ ($i=0,\dots,\theta(m)+1$) with $\deg(S_{0})\leq k$, $\deg(d_{i}S_{i})\leq k~{}(i=1,\dots,\theta(m))$, $\deg(S_{\theta(m)+1})\leq k-2$, such that
$F=S_{0}+d_{1}S_{1}+\cdots d_{\theta(m)}S_{\theta(m)}+(1-\|x\|_{2}^{2})S_{\theta(m)+1}.$
Suppose that $S_{i}=\sum\limits_{j=1}^{s_{i}}v_{i,j}v_{i,j}^{T}$ for $v_{i,j}\in\mathbb{R}[x]_{\lceil k/2\rceil}^{\ell\times 1}$ and $d_{i}=\sigma_{i}+\sum\limits_{t=1}^{r_{i}}q_{i,t}^{T}Gq_{i,t}$ for some $\sigma_{i}\in\Sigma[x]_{3^{m-1}d_{G}}^{1}$, $q_{i,t}\in\mathbb{R}[x]_{\lceil 3^{m-1}d_{G}/2-d_{G}/2\rceil}^{m\times 1}$ Then, we know that
$d_{i}S_{i}=\sigma_{i}S_{i}+\sum\limits_{j=1}^{s_{i}}\sum\limits_{t=1}^{r_{i}}v_{i,j}q_{i,t}^{T}Gq_{i,t}v_{i,j}^{T},$
which implies that $d_{i}S_{i}\in\mbox{QM}[G]_{k+3^{m-1}d_{G}}^{\ell}$. Since $(1-\|x\|_{2}^{2})\cdot I_{\ell}\in\mbox{QM}[G]^{\ell}$, we know that $1-\|x\|_{2}^{2}\in\mbox{QM}[G]_{k_{0}}$ for some $k_{0}\in\mathbb{N}$. Then, we know that $(1-\|x\|_{2}^{2})S_{\theta(m)+1}\in\mbox{QM}[G]_{k+k_{0}}^{\ell}$. Let $C$ be sufficiently large. Then, we can drop the small terms $3^{m-1}d_{G}$, $k_{0}$, and thus $F\in\mbox{QM}[G]_{k}^{\ell}$ for $k$ satisfying (5.4). Since $d_{1},\dots,d_{\theta(m)}\in\mbox{QM}[G]_{3^{m-1}d_{G}}$ and
$K=\left\{x\in\mathbb{R}^{n}\mid\frac{d_{1}(x)}{\left\|d_{1}\right\|_{B}}\geq 0,\dots,\dots,\frac{d_{\theta(m)}(x)}{\left\|d_{\theta(m)}\right\|_{B}}\geq 0\right\},$
it follows from Theorem 2.4 that $\eta\leq(3^{m-1}d_{G}+1)(3^{m}d_{G})^{n+\theta(m)-2}$.
@@
Theorem 5.1 implies that for $k$ greater than the quantity as in (5.4), we have $F\in\mbox{QM}[G]_{k}^{\ell}$. However, the bound given in (3.6) from Theorem 3.5 is generally smaller than that in (5.4). Additionally, the proof of Theorem 5.1 builds upon Theorem 3.5.
5.2 Proof of Theorem 1.2
For a polynomial matrix $F=\sum_{\alpha\in\mathbb{N}^{n}_{d}}F_{\alpha}x^{\alpha}$ with $\deg(F)=d$, $F^{hom}$ denotes its highest-degree homogeneous terms, i.e., $F^{hom}(x)=\sum_{\alpha\in\mathbb{N}^{n},|\alpha|=d}F_{\alpha}x^{\alpha}.$ The homogenization of $F$ is denoted by $\widetilde{F}$, i.e., $\widetilde{F}(\tilde{x})=\sum_{\alpha\in\mathbb{N}^{n}_{d}}F_{\alpha}x^{\alpha}x_{0}^{d-|\alpha|}$ for $\tilde{x}:=(x_{0},x)$. Let
@@
Proof 5.4.
Suppose $\tilde{x}=(x_{0},x)\in\widetilde{K}$. If $x_{0}=0$, then we have $\widetilde{F}(\tilde{x})=F^{hom}(x)\succ 0.$ If $x_{0}\neq 0$, then
@@
which implies that $x/x_{0}\in K$. Since $F^{hom}\succ 0$ on $\mathbb{R}^{n}$, the degree of $F$ must be even, and we have
@@
Since $\widetilde{K}$ is compact, the optimal value of (5.6) is positive, i.e., $\widetilde{F}_{\min}>0$.
For the polynomial matrix $(x_{0})^{d_{0}}\widetilde{G}(\tilde{x})$, let $d_{1},\dots,d_{\theta(m)}$ be the polynomials as in Theorem 4.1. Denote
(5.7) $\mathcal{G}(\tilde{x})=-\min\left(\frac{d_{1}(\tilde{x})}{\left\|d_{1}\right\|_{B}},\ldots,\frac{d_{\theta(m)}(\tilde{x})}{\left\|d_{\theta(m)}\right\|_{B}},\frac{1-\|\tilde{x}\|_{2}^{2}}{\|1-\|\tilde{x}\|_{2}^{2}\|_{B}},\frac{\|\tilde{x}\|_{2}^{2}-1}{\|1-\|\tilde{x}\|_{2}^{2}\|_{B}},0\right).$
@@
In the following, we give the proof of Theorem 1.2, which provides a polynomial bound on the complexity of matrix Putinar–Vasilescu’s Positivstellensätz.
@@
In the above, $\kappa$, $\eta$ are respectively the Łojasiewicz constant and exponent in (5.8), $C>0$ is a constant independent of $F$, $G$ and $\widetilde{F}_{\min}$ is the optimal value of (5.6). Furthermore, the Łojasiewicz exponent can be estimated as
$\eta\leq(2\cdot 3^{m-1}\lceil d_{G}/2\rceil+1)(2\cdot 3^{m}\lceil d_{G}/2\rceil)^{n+\theta(m)-1}.$
Proof 5.6.
By Lemma 5.3, we have that $\widetilde{F}\succeq\widetilde{F}_{\min}\cdot I_{\ell}\succ 0$ on $\widetilde{K}$. Note that $\mbox{QM}[1-\|\tilde{x}\|^{2}]^{\ell}$ is Archimedean, and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\widetilde{K}}$ can be equivalently described by
$d_{1}(\tilde{x})\geq 0,\dots,d_{\theta(m)}(\tilde{x})\geq 0,1-\|\tilde{x}\|_{2}^{2}\geq 0,\|\tilde{x}\|_{2}^{2}-1\geq 0.$
@@
Suppose that $S=V^{T}V$ for some polynomial matrix $V$. Then, there exist polynomial matrices $V^{(1)},V^{(2)}$ with $2\deg(V^{(1)})\leq\deg(S)$, $2\deg(V^{(2)})+2\leq\deg(S)$ such that
@@
It holds that
$\begin{array}[]{ll}S\left(\frac{(1,x)}{\sqrt{1+\|x\|_{2}^{2}}}\right)=\frac{(V^{(1)})^{T}V^{(1)}+(1+\|x\|_{2}^{2})(V^{(2)})^{T}V^{(2)}}{\sqrt{1+\|x\|_{2}^{2}}^{\deg(S)}}+\frac{(V^{(2)})^{T}V^{(1)}+(V^{(1)})^{T}V^{(2)}}{\sqrt{1+\|x\|_{2}^{2}}^{\deg(S)-1}}.\end{array}$
Similarly, there exist polynomial matrices $P_{i}^{(1)},P_{i}^{(2)}$ with $\deg(P_{i}^{(1)})\leq\deg(P_{i})$, $\deg(P_{i}^{(2)})+1\leq\deg(P_{i})$ satisfying
@@
$\begin{array}[]{l}\sum\limits_{i=1}^{t}P_{i}\left(\frac{(1,x)}{\sqrt{1+\|x\|_{2}^{2}}}\right)^{T}\frac{G(x)}{\sqrt{1+\|x\|_{2}^{2}}^{d_{0}+d_{G}}}P_{i}\left(\frac{(1,x)}{\sqrt{1+\|x\|_{2}^{2}}}\right)\\ \quad\quad\quad\quad\quad\quad=\sum_{i=1}^{t}\frac{(P_{i}^{(1)})^{T}G(x)P_{i}^{(1)}+(1+\|x\|_{2}^{2})(P_{i}^{(2)})^{T}G(x)P_{i}^{(2)}}{\sqrt{1+\|x\|_{2}^{2}}^{d_{0}+d_{G}+2\deg(P_{i})}}+\frac{(P_{i}^{(2)})^{T}G(x)P_{i}^{(1)}+(P_{i}^{(1)})^{T}G(x)P_{i}^{(2)}}{\sqrt{1+\|x\|_{2}^{2}}^{d_{0}+d_{G}+2\deg(P_{i})-1}}.\\ \end{array}$
Combining these results with equation (5.10), we know that for $k$ satisfying (5.9), the following holds:
@@
for some $Q\in\mbox{QM}[G]^{\ell}$ with $\deg(Q)\leq 2k+d$ and a symmetric polynomial matrix $H$. Since $\sqrt{1+\|x\|_{2}^{2}}$ is not a rational function, we conclude that $(1+\|x\|_{2}^{2})^{k}F=Q$, which implies $(1+\|x\|_{2}^{2})^{k}F\in\mbox{QM}[G]^{\ell}_{2k+d}$.
Since $d_{1},\dots,d_{\theta(m)}\in\mbox{QM}[G]_{2\cdot 3^{m-1}\lceil d_{G}/2\rceil}$ and
$\widetilde{K}=\left\{\tilde{x}\in\mathbb{R}^{n+1}\mid\frac{d_{1}(\tilde{x})}{\left\|d_{1}\right\|_{B}}\geq 0,\dots,\frac{d_{\theta(m)}(\tilde{x})}{\left\|d_{\theta(m)}\right\|_{B}}\geq 0,1-\|\tilde{x}\|_{2}^{2}=0\right\},$
it follows from Theorem 2.4 that $\eta\leq(2\cdot 3^{m-1}\lceil d_{G}/2\rceil+1)(2\cdot 3^{m}\lceil d_{G}/2\rceil)^{n+\theta(m)-1}$.
Proof of Corollary 1.3. Note that $F\succeq 0$ on $K$, and we have
$d+2\geq\deg(F+\varepsilon\cdot(1+\|x\|_{2}^{2})^{\lceil\frac{d+1}{2}\rceil}I_{\ell})=2\lceil\frac{d+1}{2}\rceil>d.$
Hence, for $\tilde{x}\in\widetilde{K}$, the homogenization of $F+\varepsilon\cdot(1+\|x\|_{2}^{2})^{\lceil\frac{d+1}{2}\rceil}I_{\ell}$ satisfies
$\begin{array}[]{ll}x_{0}^{2\lceil\frac{d+1}{2}\rceil-d}\widetilde{F}+\varepsilon\cdot(x_{0}^{2}+\|x\|_{2}^{2})^{\lceil\frac{d+1}{2}\rceil}I_{\ell}\succeq\varepsilon\cdot(x_{0}^{2}+\|x\|_{2}^{2})^{\lceil\frac{d+1}{2}\rceil}I_{\ell}\succeq\varepsilon\cdot I_{\ell}.\\ \end{array}$
The conclusion follows from Theorem 5.5.
5.3 Convergence rate of matrix SOS relaxations